# Shortcut to Chemically Accurate Quantum Computing via Density-based Basis-set Correction

### v1

@@ I.2 Strategy 2: Self-consistent basis-set correction
where $\bar{v}^{\mathcal{B}}(\textbf{r})$ is the derivative of the correlation basis-set correction energy with respect to the density
$\bar{v}^{\mathcal{B}}[n^{\mathcal{B}}](\textbf{r})=\frac{\delta\bar{E}^{\mathcal{B}}[n^{\mathcal{B}}]}{\delta n^{\mathcal{B}}(\textbf{r})},$ (7)
@@
and $\hat{n}^{\mathcal{B}}(\textbf{r})$ is the density operator projected in the basis set $\mathcal{B}$. This potential, and therefore the Hamiltonian, is iteratively updated from the density $n_{i}^{\mathcal{B}}$ of the wave-function ansatz solution of the $i^{\text{th}}$ iteration. The communication of the updated Hamiltonian from the CPU to the QPU/GPU is done by using standard FCIDUMP files to communicate the one- and two-electron integrals to the QPU/GPU software in order to construct the fermion operators. TREXIO files [27] are also useful to communicate a wider range of relevant information. The convergence criterion is reached when the difference between the last two iterations energies is less than 0.1 mHa. In practice, this is fulfilled after two iterations, and the self-consistent basis-set corrected energy $E_{\text{SC}}^{\mathcal{B}}$ is computed from the energy $E_{2}^{\mathcal{B}}$ coming out from the quantum solver as follows
$E_{\text{SC}}^{\mathcal{B}}=E_{2}^{\mathcal{B}}+\bar{E}^{\mathcal{B}}[n^{\mathcal{B}}]-\int\,\bar{v}^{\mathcal{B}}[n^{\mathcal{B}}](\textbf{r})n^{\mathcal{B}}(\textbf{r})\text{d}\textbf{r}.$ (8)
The final energy including the HF basis-set correction is
$E_{0}^{\mathcal{B}}=E_{\text{SC}}^{\mathcal{B}}+\Delta E_{2}^{\mathcal{B},\text{HF}},$ (9)
@@
where the HF basis-set correction is identical to the one used in Strategy 1, i.e. $\Delta E_{2}^{\mathcal{B},\text{HF}}=\Delta E_{1}^{\mathcal{B},\text{HF}}$. Additional details can be found in the SI (see Section III).
Figure 2: Ground-state errors with respect to the extrapolated CBS limit. Outside the green box: ADAPT-VQE calculations. Inside the green box: FCI calculations from a classical computer. Dot markers correspond to energies without basis-set corrections. Cross markers correspond to energies corrected with Eq. 1. On the x-axis, the labels $\mathcal{B}+$ symbolize basis-set corrected values. The blue box corresponds to a range of 1.6 mHa around the extrapolated CBS limit. The errors for each ADAPT-VQE computations with respect to the FCI energy for a given basis set are reported in the SI.
I.3 Implementation: GPU-accelerated QPU emulation and shift of the DBBSC correction to classical resources
@@
The DBBSC method does not increase the qubit count as the basis-set correction is performed classically using a density functional and a HF correction. In practice, such computations present no advantage to be performed at the QC level since they would consume a large number of qubits without accuracy benefit over its classical counterpart. Therefore a hybrid QC(ansatz)/DBBSC approach is highly preferable. Since it does not require much computational resources, ADAPT-VQE/DBBSC computations can be easily performed on a GPU-accelerated quantum emulator, the GPUs replacing the QPU. The DBBSC correction can be then shifted to the unused CPUs to maximize the computational efficiency while the computationally challenging ADAPT-VQE algorithm is fully offloaded on GPUs for optimal time-to-solution performances. Of course, the proposed extended hybrid ansatz schemes are also fully tractable on a real quantum computer, as the DBBSC correction can be entirely shifted to classical resources harnessing GPUs/CPUs to preserve the qubit count while the core ansatz is maintained on the QPU.
I.4 System-adapted basis-set generation
@@
Within the DBBSC strategy, the natural choice for basis sets is the Dunning correlation-consistent basis-set family that offers a path to a fast CBS convergence, as demonstrated in initial DBBSC classical quantum-chemistry studies [8, 7, 9, 10, 11, 12, 13, 14, 15, 16]. For quantum computing, such basis sets are unfortunately usually out of reach of quantum hardware/emulators as they are associated to very large qubit counts. In the present section and based on these large Dunning basis sets, we address the task of generating atomic-orbital (AO) basis sets under a basis-size budget, with controllable accuracy [28]. We propose a mathematical formulation of this goal in terms of a constraint optimization problem, as follows.
@@
:=$\mathcal{B}:=\{\chi_{\mu}\}_{1\leq\mu\leq N_{\text{bas}}}=\bigcup_{a=1}^{N_{\text{atm}}}\{\chi^{a}_{\mu}(\textbf{r}-\textbf{r}_{a})\}_{1\leq\mu\leq N^{a}_{\text{bas}}}$
denote a basis of atom-centered GTOs. It is assumed that an one-electron density, denoted by $n_{0}^{\mathcal{B}}$, is available after a converged ground-state energy minimization procedure (e.g. HF or CISD). Our main purpose is to a posteriori extract subsets of the given “large” basis set $\mathcal{B}$, denoted by $\mathcal{B}_{I}:=\{\chi_{\mu}\}_{\mu\in I}\subseteq\mathcal{B}$ for any index subset $I\subseteq\{1,\ldots,N_{\text{bas}}\}$, achieving (i) a target size, and (ii) minimal accuracy loss on the given density $n_{0}^{\mathcal{B}}$, used as a reference. In the present computing scheme, since the DBBSC method uses the cc-pV5Z basis set for the HF basis-set correction, we choose $\mathcal{B}$=cc-pV5Z. In practice, this choice is motivated by the fact that such a quintuple-zeta basis fairly approaches chemical accuracy compared to quasi-exact all-electron numeric atom-centered orbital computations [29]. Hence, given a target basis size $M$ smaller than $N_{\text{bas}}$, we seek the optimal index subset, denoted by $I_{M}$, that minimizes the best approximation error of the density $n_{0}^{\mathcal{B}}$ over the set of $\mathcal{B}_{I}$-representable one-electron densities, for any $I\subseteq\{1,\ldots,N_{\text{bas}}\}$ with $|I|=M$, where ||$\cdot$|| ⋅ | denotes the cardinal of a set, or, in other words, we solve an optimization problem under constraints:
@@
with $D=CC^{\top}$ the density matrix and $C$ the orbital coefficients in the AO basis. We plan to achieve this by eliminating linear dependencies present in the AO-product set [21]. The pivoted Cholesky decomposition (PCD) of a matrix [30] is an algebraic tool for eliminating linear dependencies occurring between matrix rows (resp. columns), which may be interpreted as an iterative greedy procedure for discarding elements that do not contribute to the full row (resp. column) space, up to an orthogonal projection error tolerance. PCD has been previously applied to the auxiliary basis-set generation for density fitting [19, 20, 21]. In the present work, we employ PCD within a new scheme, named system-adapted basis-set (SABS) generation, for solving the problem in Eq. (10).
$T_{pqrs}:=\sum_{\mu,\nu,\kappa,\lambda=1}^{N_{\text{bas}}}P_{p\mu}P_{q\nu}D_{\mu\nu}\langle\chi_{\mu}\chi_{\nu},\chi_{\kappa}\chi_{\lambda}\rangle D_{\kappa\lambda}P_{r\kappa}P_{s\lambda},$
and fold pairwise its first two and its last two dimensions, in order to form the $N_{\text{prod}}\times N_{\text{prod}}$ Gram matrix of weighted AO products, denoted by $G$, with $N_{\text{prod}}:=(N_{\text{orb}})^{2}$. As a last pre-processing step, we discard rows and columns of $G$ corresponding to products made of components not centered on the same atom and denote the resulting submatrix $A$. Now, PCD is applied to $A$, using the machine epsilon as a tolerance threshold, yielding an index set sorted in pivot-descending order. Given a target $M$, we define the selected AO-product index set, denoted by $S^{\text{Chol}}_{M}$, as the $M$-first pivot indices. Lastly, we recover the underlying AO index set
$J_{M}:=\bigcup_{(p_{1},p_{2})\in S^{\text{Chol}}_{M}}\{p_{1},p_{2}\}$
Note that our generation scheme directly fixes the size $M$ of the selected products, i.e. $|S_{M}^{\text{Chol}}|=M$. The actual size of the AO basis set, equal to $|I_{M}|$, is only implicitly controlled during our procedure. In practice, as numerical results show, $|I_{M}|$ is very close to $M$ for s- and p-type basis sets, while it remains the same order as $M$ for higher orbital types. Overall, the GTOs SABS generation approach is extremely fast and offers access to compact basis sets, specifically adapted to a given system and user-defined qubit budget. Examples of SABS generation can be found in the SI (see Section VI.C.).
I.5 Computational Details
In the present study, to perform the ADAPT-VQE computations, we use the Qubit-Excitation-Based pool of operators, which is considered a standard [31]. Additional details about the ADAPT-VQE methodology can be found in the SI (see section II). ADAPT-VQE computations were performed using the Hyperion-1 GPU-accelerated state-vector sparse emulator [32] up to 32 qubits. Hyperion-1 uses classical computing systems and is grounded on an efficient multi-GPU (GPU=Graphics Processing Unit) ensemble of fast custom sparse linear algebra libraries accelerating Hyperion-1’s exact/noiseless simulations. In this paper, computations were performed on NVIDIA DGX A100 nodes (8x 80GB A100 GPUs per node) and NVIDIA DGX H100 nodes (8x 80GB A100 GPUs per node). QC calculations being strongly memory-dependent, a single GPU can carry out 20 qubits ADAPT-VQE simulation depending on the nature (i.e. Hamiltonian sparsity) of the system whereas a single node (8 GPUs) can handle up to 28 qubits. Multi-node computations are required beyond such a qubit count. Further details about Hyperion-1 and its full capabilities will be given in a forthcoming publication. Convergence for ADAPT-VQE computation were set to $10^{-6}$ Hartree for all computations. Most computations started from an HF initial state. For selected ones (indicated in the text and Tables), we started the ADAPT-VQE procedure from a non-converged perturbatively selected configuration interaction CIPSI [33] initial state ($10^{-2}$ Hartree) to save some computational time within Hyperion-1. This reflects a commonly adopted strategy where a multi-determinant initial state is employed instead of a single HF determinant to increase the ground-state support of the initial state. The Quantum State Preparation of such classically-derived CIPSI wave-functions has been studied in the context of VQE [34] and QPE [35]. Also, we report in Table 1, the walltime required for several ground-state energy calculations. As one can see, the walltime is not only a function of the number of qubits: sparsity is important. For example, the H12 Hamiltonian is sparser than the one of the water molecule resulting in faster convergence. Beside the increased computing power of H100 GPUs leading to improved time-to-solution, our results also highlight the importance of fast node-to-node interconnects when performing large scale quantum emulation. Indeed, the benefit of H100 over A100 is striking for the largest 32 qubits simulations on 16 nodes where a factor 3 improvement was observed on the DGX H100 systems. Such speedup is therefore also partially related to higher node-to-node bandwidth observed on DGX H100 versus A100 systems.
@@ I.5 Computational Details
Molecule $N_{\text{adapt}}$ $N_{\text{qubits}}$ $N_{\text{gpus}}$ A100 walltime [min] H100 walltime [min]
@@ II Results
We computed the ground-state energies, dissociation curves, and dipole moments for the N2, H2O, LiH, and H2 molecules using both basis-set correction strategies. For all systems except H2, the 1s molecular orbitals were frozen, i.e. we use the frozen-core approximation. The error in each calculation is quantified as the deviation from the CBS limit, which itself is determined using a two-point extrapolation scheme, based on the cc-pVQZ and cc-pV5Z basis sets [36]. For the dipole-moment calculations we employed only Strategy 2, using an expectation value over the dipole-moment operator. The choice of this approach over a finite difference of energies is due to the lack of consistency guarantees, which arises from potential discrepancies in the treatment of the terms in the difference, thus leading to numerical errors [37]. The molecules selected for this study are commonly used in benchmarking quantum algorithms for chemistry due to their computational feasibility and the correlation strength they exhibit, which is representative of typical chemistry applications. Notably, these molecules have also been included in the systems used for benchmarking the DBBSC method in multiple studies [8, 7, 9, 10, 11, 12, 13, 14, 15, 16].
The classical computations, including basis-set corrections and the calculation of reference energies and dipole moments, were carried out using Quantum Package 2.0 [38]. In this software, the FCI energies are approximated by the energy from a truncated configuration-interaction wave function, specifically CIPSI [39], to which a second-order perturbation theory correction (PT2) is added. Given the demonstrated nearly-FCI quality of this approximation for the systems we studied, we simplify our terminology by referring to this approach as FCI rather than the more detailed CIPSI+PT2.
In this paper, we applied the basis-set corrections to the ADAPT-VQE algorithm implemented in our Hyperion-1 emulation software. Detailed technical specification are provided in the SI. Unless otherwise specified, the initial ansatz are HF wave functions. In instances where the ADAPT-VQE required an unreasonable amount of time to achieve convergence, we employed a cost-effective multi-determinant wave function as a starting point. In such cases, the number of determinants used are detailed in the notes of Table 2.

### v2

@@ I.1.2 Strategy 2: Self-consistent basis-set correction
where $\bar{v}^{\mathcal{B}}[n](\textbf{r})=\delta\bar{E}^{\mathcal{B}}[n]/\delta n(\textbf{r})$ is the derivative of the basis-set correlation correction with respect to the density, and $\hat{n}(\textbf{r})$ is the density operator.
@@
The idea is now to solve iteratively Eq. (9) on a quantum computer. The potential $\bar{v}^{\mathcal{B}}[n](\textbf{r})$, and therefore the Hamiltonian $\hat{\tilde{H}}^{\cal B}[n]$, is iteratively updated with the density $n_{i}^{\mathcal{B}}$ of the wave-function ansatz solution $\Psi^{{\cal B}}_{i}$ of the $i^{\text{th}}$ iteration. The convergence criterion is reached when the difference between the last two iterated energy eigenvalues ${\cal E}_{i}^{\cal B}$ is less than 0.1 mHa. In practice, this is fulfilled after two iterations.
Finally, the self-consistent basis-set corrected energy $E_{2}^{\cal B}$ can be directly computed from the last energy eigenvalue ${\cal E}^{\cal B}$ and density $n^{\cal B}=n_{\Psi^{\cal B}}$ coming out from the quantum solver as follows
$E_{2}^{\mathcal{B}}={\cal E}^{\cal B}+\bar{E}^{\mathcal{B}}[n^{\cal B}]-\int_{\mathbb{R}^{3}}\bar{v}^{\mathcal{B}}[n^{\cal B}](\textbf{r})n^{\cal B}(\textbf{r})\text{d}\textbf{r}+\Delta E^{\mathcal{B}}_{\text{HF}},$ (12)
@@
where the HF basis-set correction $\Delta E^{\mathcal{B}}_{\text{HF}}$ is identical to the one used in Strategy 1.
I.1.3 GPU-accelerated QPU emulation and shift of the basis-set correction to classical resources
@@
The DBBSC method does not increase the qubit count as the basis-set correction is performed classically using a density functional and a HF correction. In practice, such computations present no advantage to be performed at the QC level since they would consume a large number of qubits without accuracy benefit over its classical counterpart. Therefore a hybrid QC wave-function/classical DBBSC approach is highly preferable. In this work, VQE/DBBSC computations are performed classically on GPU-accelerated computing nodes, the GPUs replacing the quantum processing unit (QPU) through the use of a classical quantum emulator. The basis-set correction contributions can be then shifted to the unused CPUs to maximize the computational efficiency while the computationally challenging VQE algorithm is fully offloaded on GPUs for optimal time-to-solution performances. Of course, the proposed hybrid schemes are also fully tractable on a real quantum computer, as the basis-set correction contributions can be entirely shifted to classical resources harnessing GPUs/CPUs to preserve the qubit count while the core wave-function ansatz is maintained on the QPU.
I.2 System-adapted basis-set generation
@@
Within the DBBSC method, the natural choice for basis sets is the Dunning correlation-consistent basis-set family that offers a path to a reliable convergence to the CBS limit, as demonstrated in initial DBBSC classical quantum-chemistry studies [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. For QC calculations, such basis sets are unfortunately usually out of reach of quantum hardware/emulators as they are associated with very large qubit counts. In the present section, we address the task of generating atomic-orbital (AO) basis sets under a basis-size budget, with controllable accuracy [23]. We propose a mathematical formulation of this goal in terms of a constraint optimization problem, as follows.
@@
=$\mathcal{B}=\{\chi_{\mu}\}_{1\leq\mu\leq N_{\text{bas}}}=\bigcup_{a=1}^{N_{\text{atm}}}\{\chi^{a}_{\mu}(\textbf{r}-\textbf{r}_{a})\}_{1\leq\mu\leq N^{a}_{\text{bas}}}$
denote a basis set of atom-centered GTOs. It is assumed that a one-electron density, denoted by $n_{0}^{\mathcal{B}}$, is available after a converged ground-state energy minimization procedure (e.g., HF or CI). Our main purpose is to a posteriori extract subsets of the given “large” basis set $\mathcal{B}$, denoted by $\mathcal{B}_{I}=\{\chi_{\mu}\}_{\mu\in I}\subseteq\mathcal{B}$ for any index subset $I\subseteq\{1,\ldots,N_{\text{bas}}\}$, achieving (i) a target size, and (ii) minimal accuracy loss on the given density $n_{0}^{\mathcal{B}}$ used as a reference. In the present computing scheme, since the DBBSC method uses the cc-pV5Z basis set for the HF basis-set correction, we choose $\mathcal{B}$ = cc-pV5Z. In practice, this choice is motivated by the fact that such a quintuple-zeta basis set fairly approaches chemical accuracy compared to quasi-exact all-electron numeric atom-centered orbital computations [24]. Hence, given a target basis size $M$ smaller than $N_{\text{bas}}$, we seek the optimal index subset, denoted by $I_{M}$, that minimizes the best approximation error of the density $n_{0}^{\mathcal{B}}$ over the set of $\mathcal{B}_{I}$-representable one-electron densities, for any $I\subseteq\{1,\ldots,N_{\text{bas}}\}$ with $|I|=M$, where ||$\cdot$|| ⋅ | denotes the cardinal of a set, or, in other words, we solve an optimization problem under constraints:
@@
where $D=CC^{\top}$ is the density matrix and $C$ is the $N_{\text{bas}}\times N_{\text{occ}}$ matrix of the coefficients in the AO basis of the $N_{\text{occ}}$ occupied molecular orbitals. We plan to achieve this by eliminating linear dependencies present in the AO-product set [25]. The pivoted Cholesky decomposition (PCD) of a matrix [26] is an algebraic tool for eliminating linear dependencies occurring between matrix rows (resp. columns), which may be interpreted as an iterative greedy procedure for discarding elements that do not contribute to the full row (resp. column) space, up to an orthogonal projection error tolerance. PCD has been previously applied to the auxiliary basis-set generation for density fitting [27, 28, 25]. In the present work, we employ PCD within a new scheme, named system-adapted basis-set (SABS) generation, for solving the problem in Eq. (13).
@@
Molecule / Basis set $N_{\text{adapt}}$ $N_{\text{qubits}}$ $N_{\text{gpus}}$ A100 walltime [min] H100 walltime [min]
@@
$T_{pqrs}=\sum_{\mu,\nu,\kappa,\lambda=1}^{N_{\text{bas}}}P_{p\mu}P_{q\nu}D_{\mu\nu}\,\langle\chi_{\mu}\chi_{\nu},\chi_{\kappa}\chi_{\lambda}\rangle_{\text{c}}\,D_{\kappa\lambda}P_{r\kappa}P_{s\lambda},$
and fold pairwise its first two and its last two dimensions, in order to form the $N_{\text{prod}}\times N_{\text{prod}}$ Gram matrix of weighted AO products, denoted by $G$, with $N_{\text{prod}}=(N_{\text{orb}})^{2}$. As a last pre-processing step, we discard rows and columns of $G$ corresponding to products made of components not centered on the same atom and denote the resulting submatrix $A$. Now, PCD is applied to $A$, using the machine epsilon as a tolerance threshold, yielding an index set sorted in pivot-descending order. Given a target $M$, we define the selected AO-product index set, denoted by $S^{\text{Chol}}_{M}$, as the $M$-first pivot indices. Lastly, we recover the underlying AO index set
$J_{M}=\bigcup_{(p_{1},p_{2})\in S^{\text{Chol}}_{M}}\{p_{1},p_{2}\},$
Note that our generation scheme directly fixes the size $M$ of the selected products, i.e. $|S_{M}^{\text{Chol}}|=M$. The actual size of the AO basis set, equal to $|I_{M}|$, is only implicitly controlled during our procedure. In practice, as numerical results show, $|I_{M}|$ is very close to $M$ for s- and p-type basis sets, while it remains the same order as $M$ for higher angular-momentum orbital types. Overall, the SABS generation approach is extremely fast and offers access to compact basis sets, specifically adapted to a given system and user-defined qubit budget. Examples of SABS generation can be found in the SI.
We note that the adaptation and generation of basis sets to the molecular geometry has been the subject of several publications exploring other strategies. Among them, we can cite some recent works [29, 30], but also some other research in the context of quantum computing focusing on the need to limit the qubit requirements through basis-set reoptimization such as Refs. [31, 32]. Such reoptimizations cannot be performed on the fly and obtaining such improvements can take a large amount of time when dealing with large basis sets. Alternatively, Kottman and Aspuru [33] proposed a basis-set-free approach through an adaptive representation using pair-natural orbitals which was tested up to 22 qubits.
I.3 Computational Details
In the present study, we perform ADAPT-VQE computations [34] using the Qubit-Excitation-Based pool of operators, which is considered a standard [35]. Additional details about the ADAPT-VQE methodology can be found in the SI. The communication of the Hamiltonian from the CPU to the QPU/GPU is done by using standard FCIDUMP files to communicate the one- and two-electron integrals to the QPU/GPU software in order to construct the fermion operators. TREXIO files [36] are also useful to communicate a wider range of relevant information.
ADAPT-VQE computations were performed using the Hyperion-1 GPU-accelerated state-vector sparse emulator [37] up to 32 qubits. Hyperion-1 uses classical computing systems and is grounded on an efficient multi-GPU ensemble of fast custom sparse linear-algebra libraries accelerating Hyperion-1’s exact/noiseless simulations. In this paper, computations were performed on NVIDIA DGX A100 nodes (8x 80GB A100 GPUs per node) and NVIDIA DGX H100 nodes (8x 80GB A100 GPUs per node). QC calculations being strongly memory-dependent, a single GPU can carry out a 20-qubit ADAPT-VQE simulation depending on the nature (i.e. Hamiltonian sparsity) of the system whereas a single node (8 GPUs) can handle up to 28 qubits. Multi-node computations are required beyond such a qubit count. Further details about Hyperion-1 and its full capabilities will be given in a forthcoming publication. Convergence for all ADAPT-VQE computations were set to $10^{-6}$ Ha. Most computations started from a HF initial state. For selected ones (indicated in the text and Tables), we started the ADAPT-VQE procedure from a rough configuration-interaction perturbatively-selected-iteratively (CIPSI) [38] initial state (converged to only $10^{-2}$ Ha) to save some computational time within Hyperion-1. This reflects a commonly adopted strategy where a multi-determinant initial state is employed instead of a single HF determinant to increase the ground-state support in the initial state. The Quantum State Preparation of such classically-derived CIPSI wave functions has been studied in the context of VQE [39] and QPE [40]. Also, we report in Table 1, the walltime required for several ground-state energy calculations. As one can see, the walltime is not only a function of the number of qubits, sparsity is also important. For example, the Hamiltonian of the H12 molecule is sparser than the one of the water molecule resulting in a faster convergence. Beside the increased computing power of H100 GPUs leading to improved time-to-solution, our results also highlight the importance of fast node-to-node interconnects when performing large-scale quantum emulation. Indeed, the benefit of H100 over A100 is striking for the largest 32 qubits simulations on 16 nodes where an improvement of a factor 3 was observed on the DGX H100 systems. Such speedup is therefore also partially related to higher node-to-node bandwidth observed on DGX H100 versus A100 systems.
Figure 2: Ground-state energy errors with respect to the extrapolated CBS limit. Outside the green box: ADAPT-VQE calculations. Inside the green box: FCI calculations from a classical computer. Dot markers correspond to energies without basis-set corrections. Cross markers correspond to energies corrected with Strategy 1 [Eq. (6)]. On the x-axis, the labels $\mathcal{B}+$ symbolize basis-set corrected values. The blue box corresponds to a range of 1.6 mHa around the extrapolated CBS limit. The errors for each ADAPT-VQE computations with respect to the FCI energy for a given basis set are reported in the SI.
@@ II Results
We computed the ground-state energies, dissociation curves, and dipole moments for the N2, H2O, LiH, and H2 molecules using both basis-set correction strategies. For all systems except H2, the 1s molecular orbitals were frozen, i.e. we use the frozen-core approximation. Correspondingly, we use the frozen-core version of the DBBSC method as defined in Ref. [8], in which in particular the contribution to the density coming from the core electrons is neglected in the basis-set correlation density functional. The error in each calculation is quantified as the deviation from the CBS limit, which itself is determined using a two-point extrapolation scheme based on the cc-pVQZ and cc-pV5Z basis sets [21]. For the dipole-moment calculations we employed only Strategy 2, using an expectation value over the dipole-moment operator, avoiding a finite-difference approach prone to numerical errors [41].
The classical computations, including basis-set corrections and the calculation of reference energies and dipole moments, were carried out using Quantum Package 2.0 [42]. In this software, the FCI energies are approximated by the energy from a CIPSI wave function to which a second-order perturbation-theory (PT2) correction is added. Given the demonstrated nearly-FCI quality of this approximation for the systems we studied, we simplify our terminology by referring to this approach as simply FCI rather than CIPSI+PT2.
The SABS are labeled VXZ-Y where X is respectively D, T, Q, 5, or 6 for cc-pVDZ, cc-pVTZ, up to cc-pV6Z basis sets, and Y is the target size. We report in the SI the classically computed energies for the N2, H2O, LiH, and H2 molecules.